# Non-local, diamagnetic electromagnetic effects in long MITLs

### v1

@@ Acknowledgements.
The authors acknowledge useful feedback and discussions with Greg Frye, Josh Leckbee, Keith Matzen, Kate Bell, Brian Hutsel, David Sirajuddin and Christopher Jennings from SNL. This work was funded by Laboratory Directed Research and Development grant No. 229292. Sandia National Laboratories is a multimission laboratory managed and operated by National Technology & Engineering Solutions of Sandia, LLC, a wholly owned subsidiary of Honeywell International Inc., for the U.S. Department of Energy’s National Nuclear Security Administration under contract DE-NA0003525000352500035250003525. This paper describes objective technical results and analysis. Any subjective views or opinions that might be expressed in the paper do not necessarily represent the views of the U.S. Department of Energy or the United States Government.
@@ Appendix A Coaxial geometry
$E_{x}\longrightarrow\frac{E_{x}}{2\pi R},\qquad B_{\theta}=\frac{E_{r}}{c}.$ (30)

### v2

@@ Acknowledgements.
The authors acknowledge useful feedback and discussions with Greg Frye, Josh Leckbee, Keith Matzen, Kate Bell, David Sirajuddin and Christopher Jennings from SNL. Sandia National Laboratories is a multi-mission laboratory managed and operated by National Technology & Engineering Solutions of Sandia, LLC (NTESS), a wholly owned subsidiary of Honeywell International Inc., for the U.S. Department of Energy’s National Nuclear Security Administration (DOE/NNSA) under contract DE-NA0003525. This written work is authored by an employee of NTESS. The employee, not NTESS, owns the right, title and interest in and to the written work and is responsible for its contents. Any subjective views or opinions that might be expressed in the written work do not necessarily represent the views of the U.S. Government. The publisher acknowledges that the U.S. Government retains a non-exclusive, paid-up, irrevocable, world-wide license to publish or reproduce the published form of this written work or allow others to do so, for U.S. Government purposes. The DOE will provide public access to results of federally sponsored research in accordance with the DOE Public Access Plan. This work was funded by Laboratory Directed Research and Development grant No. 229292.
VIII Data availability
Data may be available upon reasonable request.
IX Author declarations
The authors have no conflicts to disclose.
@@ Appendix A Coaxial geometry
$E_{x}\longrightarrow\frac{E_{x}}{2\pi R},\qquad B_{\theta}=\frac{E_{r}}{c}.$ (A.2)